# BatteryML:An Open-source platform for Machine Learning on Battery Degradation

### 04 Dec 2023 v3

@@ 3.4 Model Selection and Implementation
BatteryML’s versatility lies in offering a spectrum of models, ensuring users can cherry-pick and tailor the most fitting analytical approach aligned with their research objectives. Anchored on the unified data format, BatteryML paves the way for integrating cutting-edge machine learning paradigms like transfer learning and multi-task learning. Such methodologies allow disparate battery datasets to synergize, sharing and transferring knowledge amongst themselves. Moreover, as we sail through an era where large-scale model architectures are blossoming, BatteryML lays a robust foundation to harness the power of these expansive models for battery research.
@@ 4.2 Battery Degradation Modeling
Table 2: Benchmark results for remaining useful life prediction. The comparison methods are split into four types, including 1) dummy regressor, a trivial baseline that uses the mean of training label as predictions; 2) linear models with features designed by domain experts; 3) traditional statistical models with QdLinear feature; 4) deep models with QdLinear feature. For models sensitive to initialization, we present the error mean across eight seeds and attach the standard deviation as subscript.
@@
Dummy regressor 398 510 419 466 331 211 460 647
“Variance” model 136 211 398 360 179 102 411 521
“Discharge” model 329 149 322 267 143 70 416 $>$1000
“Full” model 167 $>$1000 335 433 138 96 281 331
Ridge regression 116 184 $>$1000 242 169 75 490 395
PCR 90 187 653 200 197 69 $>$1000 384
PLSR 104 181 $>$1000 242 176 67 370 371
Gaussian process 154 224 $>$1000 251 204 557 559 257
XGBoost 334 799 395 720 215 84 264 324
Random forest 1689 2337 3687 53225 1922 861 3935 2110
MLP 1493 27527 4599 37081 1465 864 44810 45537
CNN 10294 228104 46575 924267 $>$1000 162116 35437 26138
LSTM 11911 21933 44329 53940 22212 977 33719 26611
Transformer 13513 36425 39111 42423 18714 935 2399 27427
@@
In the task of RUL prediction, BatteryML models predict the number of cycles until a battery’s SOH falls below a certain threshold, e.g. 80%, in comparison with the nominal capacity. These models are categorized into four types: Dummy Regressor (baseline), linear regression with expert-designed features, traditional statistical models, and deep neural network models, where the latter two types of methods are trained on QdLinear feature, i.e. the raw $Q_{d}(V_{d})$ curve.
Linear models using handcrafted features, such as the Discharge and Full model, offer relatively accurate predictions for LFP datasets. However, their performance diminishes on the MIX dataset, which features diverse aging conditions, due to the limited feature set and model capacity.
Traditional statistical models, capable of discerning non-linear patterns from low-level features such as $Q_{d}(V_{d})$ curves, employ specific modeling mechanisms such as the decision tree ensemble approach in Random Forests and the use of kernel functions in Gaussian Processes. Despite robust performance on CRUH, CRUSH, and MIX, their efficacy decreases on datasets such as MATR2 and SNL, where the number of training samples are limited. This finding indicates that these statistical models require a larger volume of training data to effectively learn and represent meaningful insights in RUL task.
@@ B.1 Remaining Useful Life Prediction
Table 2 presents the performance of various methods on the task of predicting the useful life of batteries. These methods are primarily divided into four categories. The first category, Dummy Regressor, is a baseline using the mean of the training label for prediction. The second category, including the Variance, Discharge, and Full models, comprises linear regression models based on features designed by experts in the battery domain. The Variance model considers only the variance of the QdLinear feature curve. The Discharge model takes into account characteristics during the discharge period, while the Full model also incorporates temperature and charging features. Note that the Full model does not completely encompass the features of Discharge model. For more detailed methodologies on feature construction, refer to the (Severson et al., 2019) our code implementation. The third category consists of traditional statistical models, including linear methods and tree models, which utilize the original QdLinear curves calculated by the difference of 100-th and 10-th cycles as their input feature space. The fourth category involves neural network models, with inputs being the QdLinear curves from cycle 1 to 100, minus the data from cycle 10.
@@
These models achieve significant results by fitting on low-level features, such as raw QdLinear curves. Ridge Regression displays consistent, albeit modest, performance across datasets, affirming a strong linear correlation between Qd(Vd) feature and battery cycle life. However, its underperformance on datasets like HUST, CRUSH, and MIX underscores the challenges in linear correlations under varying aging conditions. PCR and PLSR, which apply principal component analysis and project cells into different subspaces before fitting a linear regression, excel in MATR1, CRUH, and SNL datasets. This suggests that the Qd(Vd) feature exhibits a linear relationship in different subspaces with the battery life. Gaussian Process Regression (GPR) outperforms PCR and PLSR on MIX, demonstrating the importance of GPR’s non-linear fitting ability in accurately capturing the diverse degradation patterns of batteries with different electrode chemistry and operating conditions. XGBoost and Random Forest, exhibiting strong performance on CRUSH and MIX datasets, indicate that tree-based models’ non-linear capabilities are adept at effectively modeling complex battery aging patterns.
@@
Neural network models, including Multilayer Perceptron (MLP), Convolutional Neural Network (CNN), Long Short-Term Memory (LSTM), and Transformer architectures, display significant performance variability. MLP performs well on datasets such as SNL, CLO, and CRUH, while CNN shows high sensitivity to initial conditions. LSTM, exhibiting less variability, proves its robustness and superior performance on the MATR2 dataset among the neural network methods. The Transformer architecture, emerging as the most robust among deep models, achieves the best results on the CRUSH dataset across all models. These findings underscore both the potential of neural networks and the necessity for further refinement in learning methods and architectural designs to maximize their potential and possibly outperform traditional models.
In summary, there is no universally optimal method for battery modeling; all methods only excel in certain datasets and may exhibit error divergence in others. This indicates that there is considerable scope for improvement in the accuracy of early battery life prediction. Moreover, existing methods require targeted enhancements to adapt to the complex aging patterns of batteries.
@@ B.2 State of Health Estimation
This task leverages datasets from an array of sources, including CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction is comprehensive, encompassing elements such as the ratio of the current cycle’s maximum charge capacity to the nominal capacity, and the voltage of the initial cycle. We scrutinized a variety of models, including Random Forest, Linear Regression, Ridge Regression, and Neural Networks like GRU, MLP, and LSTM. Each model was run with 10 different seeds on specific data, so the values in each cell of the table represent the mean and standard deviation (std) of the RMSE results from multiple seeds. The benchmark results of the SOH estimation task are delineated in Table 4.
@@ B.3 State of Charge Estimation
Estimating the State of Charge (SOC) is contingent upon both the discharge capacity at a given moment and the SOH for the current cycle. Consequently, acquiring completely accurate SOC labels is as challenging as it is for SOH. In our methodology, we predict the SOC under the realistic workload of the battery. For this task, we employed datasets from CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction incorporates estimated current, voltage, and time information post-interpolation, alongside the charge and discharge capacity curves from preceding cycles. Labels represent the proportion of the battery’s remaining capacity at a specific point in a cycle to the current full battery capacity, multiplied by 100. The specific definition is in Formula 1. The performance of the methods of SOC prediction task is displayed in Table 5.
@@
Contrasting with the results for RUL and SOH prediction, SOC benchmark results reveal that LightGBM consistently surpasses other methodologies in the majority of tasks, thereby positioning tree models as the prevailing state-of-the-art for SOC prediction. However, all methods exhibit suboptimal performance on the RWTH dataset, suggesting that current approaches still grapple with adapting to the full spectrum of battery aging patterns. Additionally, linear models persistently outperform deep learning models, implying that both the input features and the network architecture necessitate further optimization to fully exploit the potential of deep learning models in SOC prediction tasks.
@@ C.2 Hyperparameter ablation
Figure 6 displays the hyperparameter analysis of traditional statistical methods. Figure 6a of the figure shows the performance of PLSR (Partial Least Squares Regression) with varying numbers of principal components. PLSR exhibits a trend towards overfitting with an increasing number of principal components, necessitating cross-validation to determine the optimal number of components.
Figure 6b illustrates the impact of the number of decision trees on the performance of Random Forest (RF). Random Forest shows stable effectiveness, indicating a preference for a moderate number of decision trees to balance performance and computational efficiency.
@@
Figure 7b demonstrates that Convolutional Neural Networks (CNN) require an optimal hidden dimension size, as extremes in size lead to high variance; a dimension of 16 is found to be most effective.
@@ D.1 Pipeline Interface for Efficient Replication
Almost all modules in BatteryML can be configured using the configuration file. Algorithm D.1 demonstrates an example that specifies “Variance” model on MATR1 dataset. Here we introduce these components in detail.

### 25 Mar 2024 v4

@@ 3.4 Model Development
BatteryML’s versatility lies in offering a spectrum of models, ensuring users can cherry-pick and tailor the most fitting analytical approach aligned with their research objectives. Anchored on BatteryData, BatteryML paves the way for integrating cutting-edge machine learning paradigms like transfer learning and multi-task learning into battery modeling Moreover, as we sail through an era where large-scale model architectures are blossoming, BatteryML lays a robust foundation to harness the power of these expansive models for future battery research.
@@ 4.2 Battery Degradation Modeling
Table 2: Benchmark results for remaining useful life prediction. The comparison methods are split into four types, including 1) dummy regressor, a trivial baseline that uses the mean of training label as predictions; 2) linear models with features designed by domain experts; 3) traditional statistical models with QdLinear feature; 4) deep models with QdLinear feature. For models sensitive to initialization, we present the error mean across ten seeds and attach the standard deviation as subscript.
@@
Dummy regressor 398 510 419 466 331 239 576 573
”Variance” model 136 211 398 360 179 118 506 521
”Discharge” model 329 149 322 267 143 76 $>$1000 $>$1000
”Full” model 167 $>$1000 335 433 138 93 $>$1000 331
Ridge regression 116 184 $>$1000 242 169 65 $>$1000 372
PCR 90 187 435 200 197 68 560 376
PLSR 104 181 431 242 176 60 535 383
Gaussian process 154 224 $>$1000 251 204 115 $>$1000 573
XGBoost 334 799 395 547 215 119 330 205
Random forest 1689 2337 3687 53225 1922 811 4165 1970
MLP 1493 27527 4599 37081 1465 1034 5659 45142
CNN 10294 228104 46575 924267 $>$1000 17492 54511 272101
LSTM 11911 21933 44329 53940 22212 10510 51939 2689
Transformer 13513 36425 39111 42423 18714 818 55021 27116
@@
In the task of RUL prediction, BatteryML models predict the number of cycles until a battery’s SOH falls below a certain threshold, e.g. 80%, in comparison with the nominal capacity. The performance metrics of various methods are presented in Table 2.
Linear models using handcrafted features, such as the ”Discharge” and ”Full” model, offer relatively accurate predictions for LFP battery datasets. However, their performance diminishes on the CRUSH and MIX dataset, which features diverse aging conditions, due to the limited feature set and model capacity.
Traditional statistical models, capable of discerning non-linear patterns from low-level features such as $Q_{d}(V_{d})$ curves, employ specific modeling mechanisms such as the decision tree ensemble approach in Random Forests and XGBoost. Despite robust performance on CRUH, CRUSH, and MIX, their efficacy decreases on datasets such as MATR2 and SNL, where the number of training samples are limited. This finding indicates that these statistical models require a larger volume of training data to effectively learn and represent meaningful insights in RUL task.
@@ B.1 Remaining Useful Life Prediction
Table 2 presents the performance of various methods on the task of predicting the useful life of batteries. These methods are primarily divided into four categories. The first category, Dummy Regressor, is a baseline using the mean of the training label for prediction. The second category, including the Variance, Discharge, and Full models, comprises linear regression models based on features designed by experts in the battery domain. The Variance model considers only the variance of the QdLinear feature curve. The Discharge model takes into account characteristics during the discharge period, while the Full model also incorporates temperature and charging features. Note that the Full model does not completely encompass the features of Discharge model. For more detailed methodologies on feature construction, refer to the (Severson et al., 2019; Attia et al., 2021) our code implementation. The third category consists of traditional statistical models, including linear methods and tree models, which utilize the original QdLinear curves calculated by the difference of 100-th and 10-th cycles as their input feature space. The fourth category involves neural network models, with inputs being the QdLinear curves from cycle 1 to 100, minus the data from cycle 10.
@@
These models achieve significant results by fitting on low-level features, such as raw QdLinear curves. Ridge Regression Hoerl & Kennard (2000) displays consistent, albeit modest, performance across datasets, affirming a strong linear correlation between Qd(Vd) feature and battery cycle life. However, its underperformance on datasets like HUST and CRUSH, underscores the challenges in linear correlation learning. Principal Component Regression(PCR Tipping & Bishop (1999) and artial Least Squares Regression(PLSR) Geladi & Kowalski (1986), which apply principal component analysis and project cells into different subspaces before fitting a linear regression, excel in MATR1, CRUH, and SNL datasets. This suggests that the Qd(Vd) feature exhibits a linear relationship in different subspaces with the battery life. Gaussian Process Regression (GPR) Williams & Rasmussen (2006) underperforms PCR and PLSR on all datasets, indicating that the pairwise kernel functions struggle to capture the effective degradation patterns for all kinds of batteries. XGBoost Chen & Guestrin (2016) and Random Forest Breiman (2001), exhibiting strong performance on CRUSH and MIX datasets, indicate that tree-based models’ non-linear capabilities are adept at effectively modeling complex battery aging patterns.
@@
Neural network models, including Multilayer Perceptron (MLP) Haykin (1994), Convolutional Neural Network (CNN) Krizhevsky et al. (2012), Long Short-Term Memory (LSTM) Hochreiter & Schmidhuber (1997), and Transformer Vaswani et al. (2017) architectures, display significant performance variability. MLP performs well on datasets such as SNL, CLO, and CRUH, while CNN shows high sensitivity to initial conditions. LSTM, exhibiting less variability, proves its robustness and superior performance on the MATR2 dataset among the neural network methods.
In summary, there is no universally optimal method for battery modeling; all methods only excel in certain datasets and may exhibit error divergence in others. This indicates that considerable scope exists for improvement in the accuracy of early battery life prediction. Moreover, existing methods require targeted enhancements to adapt to the complex aging patterns of batteries.
@@ B.2 State of Health Estimation
This task leverages datasets from an array of sources, including CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction is comprehensive, encompassing elements such as the ratio of the current cycle’s maximum charge capacity to the nominal capacity, and the voltage of the initial cycle. We scrutinized a variety of models, including Random Forest, Linear Regression, Ridge Regression, and Neural Networks like Gated Recurrent Unit(GRU) Cho et al. (2014), MLP, and LSTM. Each model was run with 10 different seeds on specific data, so the values in each cell of the table represent the mean and standard deviation (std) of the RMSE results from multiple seeds. The benchmark results of the SOH estimation task are delineated in Table 4.
@@ B.3 State of Charge Estimation
Estimating the State of Charge (SOC) is contingent upon both the discharge capacity at a given moment and the SOH for the current cycle. Consequently, acquiring completely accurate SOC labels is as challenging as it is for SOH. In our methodology, we predict the SOC under the realistic workload of the battery. For this task, we employed datasets from CALCE, HNEI, HUST, MATR, RWTH, SNL, and UL_PUR. Feature extraction incorporates estimated current, voltage, and time information post-interpolation, alongside the charge and discharge capacity curves from preceding cycles. Labels represent the proportion of the battery’s remaining capacity at a specific point in a cycle to the current full battery capacity, multiplied by 100. The specific definition is in Formula LABEL:eq:soc_and_soh. The performance of the methods of SOC prediction task is displayed in Table 5.
@@
Contrasting with the results for RUL and SOH prediction, SOC benchmark results reveal that LightGBM Ke et al. (2017) consistently surpasses other methodologies in the majority of tasks, thereby positioning tree models as the prevailing state-of-the-art for SOC prediction. However, all methods exhibit suboptimal performance on the RWTH dataset, suggesting that current approaches still grapple with adapting to the full spectrum of battery aging patterns. Additionally, linear models persistently outperform deep learning models, implying that both the input features and the network architecture necessitate further optimization to fully exploit the potential of deep learning models in SOC prediction tasks.
@@ C.2 Hyperparameter ablation
Figure 6 displays the hyperparameter analysis of traditional statistical methods. Figure 6a of the figure shows the performance of PLSR with varying numbers of principal components. PLSR exhibits a trend towards overfitting with an increasing number of principal components, necessitating cross-validation to determine the optimal number of components.
Figure 6b illustrates the impact of the number of decision trees on the performance of Random Forest. Random Forest shows stable effectiveness, indicating a preference for a moderate number of decision trees to balance performance and computational efficiency.
@@
Figure 7b demonstrates that CNN require an optimal hidden dimension size, as extremes in size lead to high variance; a dimension of 16 is found to be most effective.
@@ D.1 Pipeline Interface for Efficient Replication
Almost all modules in BatteryML can be configured using the configuration file. Code D.1 demonstrates an example that specifies “Variance” model on MATR1 dataset. Here we introduce these components in detail.